References MailingsReview View Help Tell me what you want to do ntain viruses. Unless you need to edit, its safer to stay in Protected View. Enable Editing Problem 7 (35 points) Henry Hawkins Industries income statement for the most recent year is given below. Sales (40,000 units) Less variable expenses Contribution margin Less fixed expenses Net income $800,000 560,000 240,000 192,000 $ 48,000 A. Compute the contribution margin per unit. B. Compute the contribution margin ratio. C. Compute the break-even point in units. D. Compute the break-even point in sales dollars. E. Compute the margin of safety in units.
media%2Ffc6%2Ffc612c64-61eb-4f7b-9f62-3c
media%2F226%2F2266ba84-d33d-4c74-bdb2-46
media%2F7ff%2F7ff09b80-3b77-4c49-8040-df
Answer #1

A)

Total Contribution margin = 240000

Number of units sold = 40000

Contribution margin per unit = Total Contribution margin/ Number of units sold

= 240000/40000 = 6

B)

Sales per unit = 800000/40000 = 20

Contribution margin ratio = [Contribution margin per unit/Sales] *100

= [6/20] *100 = 30%

C)

Break even points in units = Fixed expenses/ Contribution margin per unit

= 192000/6 = 32000 units

D)

Break even points in sales dollars = Break even points in units*Sale price per unit

= 32000*20 = 640000